1. Consider the reaction: 2BrF5 (g) <---> Br2 (g) + 5F2 (g)
  • If the equilibrium concentrations of BrF5, Br2 and F2 are 0.46 M, 1.2 M and 0.93 M, what is the value of Kc?
1
Expert's answer
2021-11-26T03:02:02-0500

Balanced chemical equation:


2BrF5(g) ⇌ Br2(g) + 5F2(g)


The equilibrium expression (Kc) for the reaction:



Hence


We know that Kp = Kc (RT)Δn




Kc = 394.53


Convert 250∘C to Kelvin (K): T = 250∘C + 273.15 = 523.15 K


Δn = (total number of moles of products) - (total number of moles in reactants)


Δn = (5 + 1) - (2) = 6 - 2 = 4


Δn = 4


R = 0.08206 L atm mol-1 K -1


Now, plug in all the numbers we found: Kp = Kc (RT)


Kp = 394.53 × (0.08206 × 523.15 K)4 = 1.34×10


Kp = 1.34×10